OpenStudy (anonymous):

What are the degree measures of all positive angles between 0 and 90 which satisfy the equation: \[\sin^2x + \cos^2x+\tan^2x+\cot^2x+\sec^2x+\csc^2x=31\]

OpenStudy (anonymous):

really? i would start with \[\tan^2(x)+\csc^2(x)=15\] but then i would be stuck

OpenStudy (anonymous):

so it's not possible?

OpenStudy (anonymous):

is possible, PITA tho

OpenStudy (anonymous):

Rewrite: \[\sin ^{2}(x)+\cos ^{2}(x)=1;\sec ^{2}(x)=\tan ^{2}(x)+1;\csc ^{2}(x)=\cot ^{2}(x)+1\] so \[2\tan ^{2}(x)+2\cot ^{2}(x)=28;\tan ^{2}(x)+\cot ^{2}(x)=14\]

OpenStudy (anonymous):

Also \[\sec ^{2}(x)+\csc ^{2}(x)=16\]

OpenStudy (anonymous):

That one goes: \[\frac {\sin ^{4}(x)+\cos ^{4}(x)}{\sin ^{2}(x)\cos ^{2}(x)}=14\] The other goes \[\frac {\sin ^{2}(x)+\cos ^{2}(x)}{\sin ^{2}(x)\cos ^{2}(x)}=16\]

OpenStudy (anonymous):

These are all Pythagorean and I am skipping steps, mostly algebraic

OpenStudy (anonymous):

yea i got to that step as well

OpenStudy (anonymous):

From here you get \[\sin ^{4}(x)+\cos ^{4}(x)=\frac {7}{8}\] Again by Pythagorean substitution you get \[1-2\cos ^{2}(x)+\cos ^{4}(x)=\frac {7}{8}\] You can use quadratic at this point and you should recognize a double angle formula

OpenStudy (anonymous):

\[\cos ^{2}(x)=\frac {2 \pm \sqrt{3}}{4}\]

OpenStudy (anonymous):

\[\frac {\sin ^{4}(x)+\cos ^{4}(x)}{\sin ^{2}(x)\cos ^{2}(x)}\div\frac {\sin ^{2}(x)+\cos ^{2}(x)}{\sin ^{2}(x)\cos ^{2}(x)}=7/8\]7/8 should equal to

OpenStudy (anonymous):

See the Pythagorean?

OpenStudy (anonymous):

oh i get it

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

how come sin^4 x is equal to 1 - cos^ x?

OpenStudy (anonymous):

\[\sin ^{4}(x)=\sin ^{2}(x)\sin ^{2}(x)=(1-\cos ^{2}(x))(1-\cos ^{2}(x))=1-2\cos ^{2}(x)+\cos ^{4}(x)\]

OpenStudy (anonymous):

\[1-2\cos ^{2}(x)+2\cos ^{4}(x)=\frac {7}{8}\]then shouldn't it be

OpenStudy (anonymous):

Oops forgot the coefficient of 2 on the fourth power, sorry

OpenStudy (anonymous):

thank you so much